Chemistry A scientist wants to make 6 milliliters of a 30% sulfuric acid solution. The solution is to be made from a combination of a 20% sulfuric acid solution and a 50% sulfuric acid solution. How many milliliters of each solution must be combined to make the 30% solution?
Answer by mananth(16946)

Acid
percent ---------------- quantity
Sulphuric acid type I 20.00% ---------------- x ml
sulphuric acid type II 50.00% ------ 6 - x ml
Mixture 30.00% ---------------- 6
Total 6 ml
20.00% x + 50.00% ( 6 - x ) = 30.00% * 6
20 x + 50 ( 6 - x ) = 180
20 x + 300 - 50 x = 180
20 x - 50 x = 180 - -300
-30 x = -120
/ -30
x = 4 ml 20.00% Sulphuric acid type I
2 ml 50.00% sulphuric acid type II

m.ananth@hotmail.ca